OpenStudy (anonymous):

The beginning tennis class has twice as many students as the advanced class. The intermediate class has three more students than the advanced class. How many students are in the advanced class if the total enrollment for the three tennis classes is 39?

OpenStudy (butterflydreamer):

let's let b = beginner class , a = advanced class and i = intermediate class. So from the information we have we have the following equations: "The beginning class has twice as many as many students as the advanced class" Therefore, b = 2a right? :) Then "the intermediate class has three more students than the advanced class" Therefore, i = a + 3 Then we know that the TOTAL number of students is 39. Therefore, b + a + i = 39 Substitute b = 2a and i = (a+3) into the "b + a + i = 39"

OpenStudy (butterflydreamer):